OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2).

OpenStudy (shinalcantara):

any other data given?

OpenStudy (anonymous):

y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1

OpenStudy (shinalcantara):

since you have the line with points (0, -1) and (-2, 0) then we can solve for the slope of the line Formula would be: \[m = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] substitute the formula using the points to get the slope

OpenStudy (anonymous):

ok thanks